OpenStudy (anonymous):

What is the solution set of |x – 6| = 8?

OpenStudy (anonymous):

The solution set will be all values of x which make the equation true. When a number is enclosed like this |number| it means that the absolute, or positive value is taken - ignore the sign so that |-x| = x. You want to find all values of x which satisfy |x-6| = 8. Because of the |..| you are looking for both negative and positive values: what value of x will make x-6 = -8, and what value of x will make x-6 = +8?